OpenStudy (anonymous):

If f(x) = −4x − 1, find f(−2)

OpenStudy (anonymous):

Substitute the value of x with -2 to get the answer.

OpenStudy (amorfide):

sub in -2 into x in the equation

OpenStudy (lgbasallote):

replace every x of f(x) with -2

Parth (parthkohli):

f(x) = -4x - 1 Therefore, \(\Large \color{MidnightBlue}{\Rightarrow -4(-2) - 1 }\) This can be used.

OpenStudy (anonymous):

f(-2) = -4(-2) - 1 = 8 - 1 = 7

Parth (parthkohli):

x is like a thing that is used in general